Definition, Formula and Physical interpretation of Cross Product.


1
Expert's answer
2021-11-15T16:53:03-0500

The cross product a × b is defined as a vector c that is perpendicular (orthogonal) to both a and b, with a direction given by the right-hand rule and a magnitude equal to the area of the parallelogram that the vectors span.

The cross product is defined by the formula


"\vec a\times\vec b=||\vec a||\cdot||\vec b||\sin \theta \vec n"

where:

"\u03b8"  is the angle between "\vec a"  and "\vec b"  in the plane containing them (hence, it is between 0° and 180°)

"||\vec a||" and "||\vec b||" are the magnitudes of vectors "\vec a" and "\vec b"

and "\vec n"  is a unit vector perpendicular to the plane containing  "\vec a" and "\vec b" , in the direction given by the right-hand rule .

If the vectors  "\vec a" and "\vec b"  are parallel (that is, the angle θ between them is either 0° or 180°), by the above formula, the cross product of "\vec a" and "\vec b"  is the zero vector  "\vec 0."


If "(i, j,k)" is a positively oriented orthonormal basis, the


"\vec a\times\vec b=\begin{vmatrix}
 i & j & k \\
 a_1 & a_2 & a_3 \\
 b_1 & b_2 & b_3 \\
\end{vmatrix}"

The cross product is used to describe the Lorentz force experienced by a moving electric charge "q_e:"


"\vec F=q_e(\vec E+\vec v\times \vec B)"

Since velocity "\vec v,"  force "\vec F"  and electric field "\vec E"  are all true vectors, the magnetic field "\vec B"  is a pseudovector.


The moment "\vec M"  of a force "\vec F_B"  applied at point B around point A is given as:


"\vec M_A=\vec r_{AB}\times\vec F_B"

The angular momentum "\vec L"  of a particle about a given origin is defined as:


"\vec L=\vec r\times \vec p,"

where "\vec r"  is the position vector of the particle relative to the origin, "\vec p"   is the linear momentum of the particle.